# reference.bib

### v1

@@ 1 Introduction
Hand-object interaction (HOI) involves jointly modeling hand articulation and object dynamics to generate and interpret realistic manipulation sequences [grauman2024egoexo4dunderstandingskilledhuman, kwon2021h2ohandsmanipulatingobjects, liu2024tacobenchmarkinggeneralizablebimanual, liu2024hoi4d4degocentricdataset]. This reflects one of the most pervasive human behaviors, deeply embedded in daily activities. Generating and understanding 3D HOI sequences is critical for advancing machine capabilities in human-centric applications. In augmented/virtual reality (AR/VR), realistic HOI modeling enables immersive digital experiences, allowing users to manipulate virtual objects naturally. For robotics, it provides the foundation for dexterous, feedback-driven manipulation in unstructured environments.
@@ 2 Related Work
Synthesizing realistic and diverse hand-object interactions (HOI) have gained significant attention, with numerous approaches exploring this problem under various settings. These include the creation of large-scale datasets like GigaHands [fu2025gigahands] and OAKINK2 [zhan2024oakink2], methods for affordance learning [Li_2024, zhang2025bimartunifiedapproachsynthesis, Prakash2025LatentAct, wei2025afforddexgrasp]. Others have focused on synthesizing two-hand interactions like InterHandGen [lee2024interhandgen] and HandDiffuse [lin2025handdiffuse], or on physics-aware synthesis and dexterous grasping [wang2024unigrasptransformer, wei2024grasp, zhong2025dexgrasp]. Several key works have specifically tackled text-guided or semantically rich HOI synthesis. Text2HOI [cha2024text2hoi] introduced a pioneering approach to generate 3D HOI motion from text by decomposing the task into contact prediction and motion generation using a diffusion model, but it can struggle with fine-grained control from low-level text and tends to produce short HOI interactions. HOIGPT [huang_etal_cvpr25] adopts an LLM-based sequential model to predict hand-object trajectories from text for long sequences, yet it lacks explicit affordance-guided mechanisms and cannot ensure smooth motion in-between multiple sub-sequences in complex interactions. SemGrasp [li2024semgrasp] focuses on semantic-aware grasp generation, but it discretizes motion space and is limited to static grasps. Grasp as You Say [wei2024grasp] addresses category-level language-guided grasping, but it requires hand-crafted hand-object contact templates and lacks dynamics. While these methods have advanced the field, generating fine-grained, diverse, and long-horizon hand-object interaction sequences with high-level text on unseen objects cohesively remains a significant challenge.
@@
The advent of Large Language Models (LLMs) has revolutionized natural language understanding, and their influence is gradually expanding into 3D perception and interaction tasks. Some recent works leverage pure LLM architectures to handle motion generation from language. For instance, HOIGPT [huang_etal_cvpr25] and MotionGPT [jiang2024motiongpt] adopt VQ-VAE tokenization and GPT-style architectures to achieve bidirectional generation between natural language and HOI systhesis. However, these models operate primarily on discrete motion tokens, and lack grounding in actual 3D perception or reasoning, they cannot perceive objects or environments in 3D nor reason about spatial affordances or constraints, thus limiting their ability to generate context-aware interactions. Conversely, Multimodal Large Language Models (MLLMs) have emerged as powerful tools that extend the success of LLMs into various domains, aiming to bridge the gap between language and other modalities such as images, videos, and 3D representations, enabling tasks like visual question answering like ShapeLLM [qi2024shapellm], 3D object generation like Point-E [nichol2022pointegenerating3dpoint] and CLIP-Forge [sanghi2022clipforgezeroshottexttoshapegeneration], and affordance understanding like LASO [li2024laso] and AffordanceLLM [qian2024affordancellm]. These MLLMs demonstrate strong capabilities in static 3D understanding, especially in segmenting or identifying affordances from language prompts. Other MLLM-based approaches like SeqAfford [yu2025seqaffordsequential3daffordance] and 3D-AffordanceLLM [chu20253d] further extend this to sequential or compositional affordance reasoning. However, these approaches remain limited to perception tasks and do not address the synthesis of continuous hand-object interactions or motion dynamics. In contrast, while LLM-based methods focus on generating motion from text without grounding in 3D environments, and MLLM-based methods enable 3D spatial understanding without supporting interaction generation, none of the existing approaches bridge both capabilities. Our work addresses this gap by leveraging an MLLM that is trained to jointly model language, 3D perception, and interaction dynamics. This enables our model to both understand object affordances from language and synthesize long-horizon, task-consistent hand-object interactions with unseen objects from open-vocabulary instructions. By unifying reasoning and generation, our approach extends 3D MLLMs toward fine-grained, dynamic, and open-world HOI synthesis.
@@ 3.1 Instruction Decomposition and Affordance Reasoning via 3D MLLM
Given a 3D object point cloud and a high-level instruction (e.g., "I am thirsty"), the 3D MLLM semantically grounds the instruction into object-centric affordances by bridging the gap between abstract intent and actionable object functionality. This process yields two structured outputs: (1) A spatial affordance map that identifies geometrically plausible interaction regions (e.g., highlighting the cap of a water bottle for opening and the body for grasping). (2) A temporally decomposed sub-task sequence that translates the instruction into executable atomic actions (e.g., 1) grasp bottle cap with both hands → 2) twist counterclockwise to open → 3) lift bottle to mouth with right hand").
@@
Although 3D MLLMs effectively align 3D representations with natural language, they are predominantly tailored for object-centric text generation and thus lack inherent support for dense 3D prediction tasks such as fine-grained affordance segmentation. To address this limitation, we augment the MLLM vocabulary with a dedicated segmentation token <AFF>, following the design of Lisa [lai2024lisa], thereby enabling the model to represent and reason about segmentation outputs within its language-based framework. Formally, given a point-cloud input $\mathbf{F}_{\text{obj}}$ and instruction text $\mathbf{T}_{\text{ins}}$ expressing user intent over candidate objects, the 3D MLLM jointly encodes both modalities to produce a sub-task sequence:
@@
which contains $S$ occurrences of the segmentation token <AFF>, each marking a predicted affordance region. We then collect the corresponding last-layer embeddings$\{\mathbf{h}_{\text{aff}}^{(i)}\}_{i=0}^{S-1}$and pass each through an affordance decoder:
@@
Our training pipeline comprises two sequential stages. First, we fine-tune the model on large-scale, coarse-grained, object-centric static affordance datasets [qian2024affordancellm, yu2025seqaffordsequential3daffordance], to instill strong affordance priors, thereby enhancing its generalization and open-vocabulary capabilities. Second, we fine-tune the stage 1 model on a smaller dynamic hand-object interaction dataset, which can get fine-grained affordance maps by using the voxel-based method, to generate more precise affordance Maps and task decomposition sequences. We leverage an auto-regressive cross-entropy loss $L_{\text{task}}$ to supervise sub-tasks generation, complemented by Dice loss and Binary Cross-Entropy loss $L_{\text{aff}}$ to guide affordance prediction.
@@ 3.2 Affordance-driven HOI Diffusion with Physical Refinement
During training, given a conditioning pair $[X,\textbf{C}]$, we first add noise to $X$ using the forward diffusion process:
@@
where $L_{\text{hoi\_diff}}(\hat{X}_{\theta},\textbf{C})=\mathbb{E}_{X_{0}\sim p(X_{0}),X_{t}\sim p(X_{t}|X_{0}),t\sim[1,T]}||X_{0}-X_{\theta}(X_{t},t,\textbf{C})||_{2}^{2}$ is the denoising loss of the diffusion process for HOI reconstruction. $L_{\text{hoi\_distance}}(\hat{X}_{\theta})$ donates the HOI distance map loss, which encourages precise surface contact and enhances the physical plausibility of the contact area by penalizing errors more strongly when the hand is near the object. $L_{\text{hoi\_orient}}(\hat{X}_{\theta})$ represents the HOI relative orientation loss, which aligns predicted orientations with ground truth to ensure accurate rotational poses for tasks like grasping and manipulation by modeling hand–object relative rotation, respectively. To better align the generated HOI sequences with the Interaction Prior C, we adopt the classifier-free guidance strategy:
@@
where $d(\cdot,\cdot)$ is the Euclidean distance between hand joints $(\hat{J}_{\text{lhand}},\hat{J}_{\text{rhand}})$ and closest object surfaces $(\tilde{P}_{\text{obj}}^{\text{ljoint}},\tilde{P}_{\text{obj}}^{\text{rjoint}})$.
@@
where $d(\cdot,\cdot)$ is the Euclidean distance between hand joints $(\hat{V}_{\text{lhand}},\hat{V}_{\text{rhand}})$ and closest object surfaces $(\tilde{P}_{\text{obj}}^{\text{lvert}},\tilde{P}_{\text{obj}}^{\text{rvert}})$.
@@
$l_{\text{transition}}=||\hat{V}_{\text{trans}}^{0}-V_{\text{pre}}^{T}||_{2}^{2}+||\hat{V}_{\text{trans}}^{T}-V_{\text{after}}^{0}||_{2}^{2},$ (11)
@@ 4.1 Dataset
For our experiments, we utilize two prominent hand-object interaction datasets: GRAB [GRAB:2020], which provides comprehensive full-body motion data of subjects interacting with 51 everyday objects, and ARCTIC [fan2023arctic], a large-scale dataset specializing in bi-manual interactions with articulated objects and dense 3D annotations. We preprocess both datasets with a unified pipeline, which involves initial geometric operations on the object point clouds—specifically, upsampling to support detailed affordance map generation via our inference model, followed by downsampling to ensure computational efficiency during training. Crucially, we enrich the original annotations by employing a large language model (LLM) to convert low-level HOI descriptions into open-vocabulary, intent-centric language instructions. These semantic labels enhance the expressiveness of the data and allow our model to better generalize to unseen scenarios, supporting open-world HOI synthesis with long-horizon sequences and diverse objects. For both GRAB [GRAB:2020] and ARCTIC [fan2023arctic], we follow a standard protocol by partitioning each dataset into 80% for training and 20% for unseen testing, ensuring reliable evaluation of our model’s generalization capabilities. This setup allows us to rigorously assess performance on unseen objects, motions, and interaction intents, validating the robustness of our MLLM-guided generation framework across both single-hand and bi-manual interaction scenarios.
@@ 4.2 Evaluation metric
To comprehensively evaluate the quality, diversity, realism, and physical plausibility of our generated hand-object interaction (HOI) sequences, we employ a multi-faceted set of quantitative and qualitative metrics inspired by prior works [cha2024text2hoi, huang_etal_cvpr25, liu2024geneoh, luo2024physics, tian2024gaze]. We roughly divide the evaluation indicators into three categories
  • Motion Accuracy. Mean Per-Joint Position Error (MPJPE) for hand joints and Final Object Location Error (FOL) for object placement
  • Generation realism. Fréchet Inception Distance (FID) between real and synthesized motions.
  • Diversity (Diversity) and multi-modality (MModality) scores, which measure the variability across different prompts and within outputs from the same prompt .
Lower MPJPE, FOL, and FID indicate higher accuracy and fidelity, while higher diversity and multi-modality reflect stronger generative expressiveness.
@@ 4.3 Main Results
MDM[tevet2023human] 74.92$\pm$2.25 0.62$\pm$0.02 62.37$\pm$1.56 3.28$\pm$0.10 12.77$\pm$0.45
TM2T[chuan2022tm2t] 59.27$\pm$1.19 0.46$\pm$0.06 57.41$\pm$2.30 3.60$\pm$0.07 21.28$\pm$0.82
MotionGPT[jiang2024motiongpt] 63.94$\pm$2.56 0.43$\pm$0.01 52.03$\pm$1.82 3.61$\pm$0.08 20.26$\pm$0.51
Text2HOI[cha2024text2hoi] 56.29$\pm$2.13 0.44$\pm$0.03 33.72$\pm$1.27 3.41$\pm$0.16 17.71$\pm$0.87
Ours 47.64$\pm$1.03 0.26$\pm$0.02 26.43$\pm$0.77 3.69$\pm$0.27 24.59$\pm$2.01
Unseen MDM[tevet2023human] 92.97$\pm$1.86 0.69$\pm$0.03 75.59$\pm$1.89 3.07$\pm$0.11 11.15$\pm$0.85
TM2T[chuan2022tm2t] 61.07$\pm$1.34 0.55$\pm$0.02 66.43$\pm$1.66 3.37$\pm$0.07 14.03$\pm$0.67
MotionGPT[jiang2024motiongpt] 66.26$\pm$1.99 0.51$\pm$0.01 56.49$\pm$1.98 2.852$\pm$0.07 16.36$\pm$0.53
Text2HOI[cha2024text2hoi] 60.67$\pm$1.80 0.41$\pm$0.02 36.96$\pm$0.77 1.80$\pm$0.05 10.98$\pm$0.44
Ours 51.34$\pm$0.85 0.27$\pm$0.01 28.29$\pm$0.62 3.61$\pm$0.09 19.91$\pm$0.63
@@
MDM[tevet2023human] 72.67$\pm$0.63 0.60$\pm$0.05 33.66$\pm$0.19 2.35$\pm$0.05 8.20$\pm$0.20
TM2T[chuan2022tm2t] 54.39$\pm$0.64 0.41$\pm$0.04 34.12$\pm$0.49 1.67$\pm$0.02 13.60$\pm$0.17
MotionGPT[jiang2024motiongpt] 60.17$\pm$0.72 0.41$\pm$0.03 31.58$\pm$0.46 1.89$\pm$0.02 13.23$\pm$0.09
Text2HOI[cha2024text2hoi] 52.16$\pm$0.41 0.33$\pm$0.01 23.35$\pm$0.33 2.43$\pm$0.02 11.21$\pm$0.20
Ours 45.15$\pm$0.94 0.25$\pm$0.04 19.74$\pm$0.16 2.65$\pm$0.03 15.25$\pm$1.44
Unseen MDM[tevet2023human] 86.75$\pm$1.35 0.64$\pm$0.01 41.53$\pm$1.37 1.58$\pm$0.04 7.13$\pm$0.63
TM2T[chuan2022tm2t] 55.57$\pm$1.26 0.53$\pm$0.03 37.22$\pm$0.75 1.54$\pm$0.12 11.23$\pm$0.44
MotionGPT[jiang2024motiongpt] 64.41$\pm$0.73 0.43$\pm$0.04 33.99$\pm$2.43 1.50$\pm$0.09 11.08$\pm$0.79
Text2HOI[cha2024text2hoi] 57.83$\pm$1.61 0.39$\pm$0.01 25.22$\pm$0.59 1.61$\pm$0.06 7.11$\pm$0.25
Ours 47.25$\pm$0.39 0.28$\pm$0.03 20.05$\pm$0.80 2.49$\pm$0.08 12.66$\pm$0.71

### v2

# OpenHOI: Open-World Hand-Object Interaction Synthesis with Multimodal Large Language Model

@@ 1 Introduction
Hand-object interaction (HOI) involves jointly modeling hand articulation and object dynamics to generate and interpret realistic manipulation sequences [grauman2024egoexo4dunderstandingskilledhuman, kwon2021h2ohandsmanipulatingobjects, liu2024tacobenchmarkinggeneralizablebimanual, liu2024hoi4d4degocentricdataset, ji2025immersivehumanxinteractionrealtime, wu2024thortexthumanobjectinteraction, deng2025humanobjectinteractionautomaticallydesigned]. This reflects one of the most pervasive human behaviors, deeply embedded in daily activities. Generating and understanding 3D HOI sequences is critical for advancing machine capabilities in human-centric applications. In augmented/virtual reality (AR/VR), realistic HOI modeling enables immersive digital experiences, allowing users to manipulate virtual objects naturally. For robotics, it provides the foundation for dexterous, feedback-driven manipulation in unstructured environments.
@@ 2 Related Work
Synthesizing realistic and diverse hand-object interactions (HOI) have gained significant attention, with numerous approaches exploring this problem under various settings. These include the creation of large-scale datasets like GigaHands [fu2025gigahands] and OAKINK2 [zhan2024oakink2], methods for affordance learning [Li_2024, zhang2025bimartunifiedapproachsynthesis, Prakash2025LatentAct, wei2025afforddexgrasp]. Others have focused on synthesizing two-hand interactions like InterHandGen [lee2024interhandgen] and HandDiffuse [lin2025handdiffuse], or on physics-aware synthesis and dexterous grasping [wang2024unigrasptransformer, wei2024grasp, zhong2025dexgrasp]. Several key works have specifically tackled text-guided or semantically rich HOI synthesis. Text2HOI [cha2024text2hoi] introduced a pioneering approach to generate 3D Hand-Object Interaction Sequences from text by decomposing the task into contact prediction and motion generation using a diffusion model, but it can struggle with fine-grained control from low-level text and tends to produce short HOI interaction sequences. HOIGPT [huang_etal_cvpr25] adopts an LLM-based sequential model to predict hand-object trajectories from text for long-horizon sequences, yet it lacks explicit affordance-guided mechanisms and cannot ensure smooth motion in-between multiple sub-sequences in complex interactions. SemGrasp [li2024semgrasp] focuses on semantic-aware grasp generation, but it discretizes motion space and is limited to static grasps. Grasp as You Say [wei2024grasp] addresses category-level language-guided grasping, but it requires hand-crafted hand-object contact templates and lacks dynamics. While these methods have advanced the field, generating fine-grained, diverse, and long-horizon hand-object interaction sequences with high-level text on unseen objects cohesively remains a significant challenge.
@@
The advent of Large Language Models (LLMs) has revolutionized natural language understanding, and their influence is gradually expanding into 3D perception and interaction tasks. Some recent works leverage pure LLM architectures to handle motion generation from language. For instance, HOIGPT [huang_etal_cvpr25] and MotionGPT [jiang2024motiongpt] adopt VQ-VAE tokenization and GPT-style architectures to achieve bidirectional generation between natural language and HOI systhesis. However, these models operate primarily on discrete motion tokens, and lack grounding in actual 3D perception or reasoning, they cannot perceive objects or environments in 3D nor reason about spatial affordances or constraints, thus limiting their ability to generate context-aware interactions. Conversely, Multimodal Large Language Models (MLLMs) have emerged as powerful tools that extend the success of LLMs into various domains, aiming to bridge the gap between language and other modalities such as images, videos, and 3D representations, enabling tasks like visual question answering like ShapeLLM [qi2024shapellm], 3D object generation like Point-E [nichol2022pointegenerating3dpoint] and CLIP-Forge [sanghi2022clipforgezeroshottexttoshapegeneration], and affordance understanding like LASO [li2024laso], DAG[wang2025dagunleashpotentialdiffusion]and AffordanceLLM [qian2024affordancellm]. These MLLMs demonstrate strong capabilities in static 3D understanding, especially in segmenting or identifying affordances from language prompts. Other MLLM-based approaches like SeqAfford [yu2025seqaffordsequential3daffordance] and 3D-AffordanceLLM [chu20253d] further extend this to sequential or compositional affordance reasoning. However, these approaches remain limited to perception tasks and do not address the synthesis of continuous hand-object interactions or motion dynamics. In contrast, while LLM-based methods focus on generating motion from text without grounding in 3D environments, and MLLM-based methods enable 3D spatial understanding without supporting interaction generation, none of the existing approaches bridge both capabilities. Our work addresses this gap by leveraging an MLLM that is trained to jointly model language, 3D perception, and interaction dynamics. This enables our model to both understand object affordances from language and synthesize long-horizon, task-consistent hand-object interactions with unseen objects from open-vocabulary instructions. By unifying reasoning and generation, our approach extends 3D MLLMs toward fine-grained, dynamic, and open-world HOI synthesis.
@@ 3.1 Instruction Decomposition and Affordance Reasoning via 3D MLLM
Given a 3D object point cloud and a high-level,free-form instruction (e.g., "I am thirsty"), the 3D MLLM semantically grounds the instruction into object-centric affordances by bridging the gap between abstract intent and actionable object functionality. This process yields two structured outputs: (1) A spatial affordance map that identifies geometrically plausible interaction regions (e.g., highlighting the cap of a water bottle for opening and the body for grasping). (2) A temporally decomposed sub-task sequence that translates the instruction into executable atomic actions (e.g., 1) grasp bottle cap with both hands → 2) twist counterclockwise to open → 3) lift bottle to mouth with right hand").
@@
Although 3D MLLMs effectively align 3D representations with natural language, they are predominantly tailored for object-centric text generation and thus lack inherent support for dense 3D prediction tasks such as fine-grained affordance segmentation. To address this limitation, we augment the MLLM vocabulary with a dedicated segmentation token <AFF>, following the design of Lisa [lai2024lisa], thereby enabling the model to represent and reason about segmentation outputs within its language-based framework. Formally, given a point‐cloud input $\mathbf{F}_{\text{obj}}$ and instruction text $\mathbf{T}_{\text{ins}}$ expressing user intent over candidate objects, the 3D MLLM jointly encodes both modalities to produce a sub-task sequence:
@@
[…]
@@
Our training pipeline comprises two sequential stages. First, we fine-tune the model on large-scale, coarse-grained, object-centric static affordance datasets [qian2024affordancellm, yu2025seqaffordsequential3daffordance], to instill strong affordance priors, thereby enhancing its generalization and open-vocabulary capabilities. Second, we fine-tune the stage 1 model on a smaller dynamic hand-object interaction dataset, which can get fine-grained affordance maps by using the voxel-based method, to generate more precise affordance Maps and task decomposition sequences. We leverage an auto‐regressive cross‐entropy loss $L_{\text{task}}$ to supervise sub-tasks generation, complemented by Dice loss and Binary Cross-Entropy loss $L_{\text{aff}}$ to guide affordance prediction.
@@ 3.2 Affordance-driven HOI Diffusion with Physical Refinement
During training, given a conditioning pair $X,\textbf{C}]$, we first add noise to $X$ using the forward diffusion process:
@@
[…]
@@
where $d(\cdot,\cdot)$ is the Euclidean distance between hand joints $(\hat{J}_{\text{lhand}},\hat{J}_{\text{rhand}})$ and closest affordance region $(\tilde{P}_{\text{obj}}^{\text{ljoint}},\tilde{P}_{\text{obj}}^{\text{rjoint}})$.
@@
where $d(\cdot,\cdot)$ is the Euclidean distance between hand joints $(\hat{V}_{\text{lhand}},\hat{V}_{\text{rhand}})$ and closest affordance region $(\tilde{P}_{\text{obj}}^{\text{lvert}},\tilde{P}_{\text{obj}}^{\text{rvert}})$.
@@
[…]
@@ 4.1 Dataset
For our experiments, we utilize two prominent hand-object interaction datasets: GRAB [GRAB:2020], which provides comprehensive full-body motion data of subjects interacting with 51 everyday objects, and ARCTIC [fan2023arctic], a large-scale dataset specializing in bi-manual interactions with articulated objects and dense 3D annotations. We preprocess both two datasets with a unified pipeline, which involves initial geometric operations on the object point clouds—specifically, upsampling to support detailed affordance map generation via our inference model, followed by downsampling to ensure computational efficiency during training. Crucially, we enrich the original annotations by employing a multimodal large language model (MLLM) to convert low-level HOI motion descriptions into open-vocabulary, intent-centric language instructions. These semantic labels enhance the expressiveness of the data and allow our model to better generalize to unseen scenarios, supporting open-world HOI synthesis with long-horizon sequences and diverse objects. For both GRAB [GRAB:2020] and ARCTIC [fan2023arctic], we follow a standard protocol by partitioning each dataset into 80% for training and 20% for unseen testing, ensuring reliable evaluation of our model’s generalization capabilities. This setup allows us to rigorously assess performance on unseen objects, motions, and interaction intents, validating the robustness of our MLLM-guided generation framework across both single-hand and bi-manual interaction scenarios.
@@ 4.2 Evaluation metric
To comprehensively evaluate the quality, diversity, realism, and physical plausibility of our generated hand-object interaction (HOI) sequences, we employ a multi-faceted set of quantitative and qualitative metrics inspired by prior works [cha2024text2hoi, huang_etal_cvpr25, liu2024geneoh, luo2024physics, tian2024gaze, hcfmm]. We roughly divide the evaluation indicators into three categories
  • Motion Accuracy. We evaluate geometric precision using the Mean Per-Joint Position Error (MPJPE;) computed over hand joints, and assess object placement with the Final Object Location Error (FOL;), defined as the Euclidean distance between the predicted object center and the target location at the final frame.
  • Generation Realism. We measure realism via the Fréchet Inception Distance (FID;) between real and synthesized motions in a pre-trained motion feature space, capturing distributional alignment and perceptual fidelity.
  • Diversity & Multi-modality. Diversity quantifies across-prompt variability of generated outputs, while MModality captures within-prompt variability across multiple samples. Both are computed from pairwise distances or variance statistics in the motion feature space.
Lower MPJPE, FOL, and FID indicate higher accuracy and fidelity, while higher multi-modality and closer to GT diversity reflect stronger generative expressiveness.
@@ 4.3 Main Results
[…]
@@
[…]